6x+4y=20
4x-2y=4
solving systems of equations by using the elimination method with addition or subtraction or multiplication ...

Answer by rmromero(383)

6x+4y=20
4x-2y=4 
solving systems of equations by using the elimination method with addition or subtraction or multiplication ...


Elimination Method: we are to eliminate one variable. Choose y

           6x + 4y = 20    Notice that if you add 4y and -2y, 
           4x - 2y = 4     we cant eliminate y.

Multiply 2 to 4x - 2y = 4

          6x + 4y = 20    
        2(4x - 2y = 4)

         6x + 4y = 20    
         8x - 4y = 8
        _____________
         14x     = 28         Solve for x
          14x/14 = 28/14
               x = 2

Solve y by substituting x = 2 to any of the two equations

            4x - 2y = 4 , x = 2
          4(2) - 2y = 4
             8 - 2y = 4
                -2y = -4
                  y = 2

Check if the answers x = 2 and y = 2 makes the two equation true after
substituting the value of x and y to the two equations.


Happy calculating


Rachel